# FractalCloud: A Fractal-Inspired Architecture for Efficient Large-Scale Point Cloud Processing

### v1

@@ Abstract
Three-dimensional (3D) point clouds are increasingly used in applications such as autonomous driving, robotics, and virtual reality (VR). Point-based neural networks (PNNs) have demonstrated strong performance in point cloud analysis, originally targeting small-scale inputs. However, as PNNs evolve to process large-scale point clouds with hundreds of thousands of points, all-to-all computation and global memory access in point cloud processing introduce substantial overhead, causing $O(n^{2})$ computational complexity and memory traffic where $n$ is the number of points. Existing accelerators, primarily optimized for small-scale workloads, overlook this challenge and scale poorly due to inefficient partitioning and non-parallel architectures. To address these issues, we propose FractalCloud, a fractal-inspired hardware architecture for efficient large-scale 3D point cloud processing. FractalCloud introduces two key optimizations: (1) a co-designed Fractal method for shape-aware and hardware-friendly partitioning, and (2) block-parallel point operations that decompose and parallelize all point operations. A dedicated hardware design with on-chip fractal and flexible parallelism further enables fully parallel processing within limited memory resources. Implemented in 28 nm technology as a chip layout with a core area of 1.5 $mm^{2}$, FractalCloud achieves 21.7× speedup and 27× energy reduction over state-of-the-art accelerators while maintaining network accuracy, demonstrating its scalability and efficiency for PNN inference.
11footnotetext: Corresponding Author
@@ I Introduction
With the rapid advancement of 3D sensing technologies [rani2024advancements, guo2024revisiting, yan2024spot, zhou2023sagitta], such as LiDAR [roriz2024survey] and stereo cameras [dong20231920, dong20214], point cloud has become prevalent in a wide range of edge applications, including autonomous driving [cui2021deep], robotics [kastner20203d], augmented reality (AR)[wang2018point], and virtual reality (VR)[stets2017visualization, li2024fusion]. The points in a 3D point cloud are inherently unordered and irregularly distributed, making it challenging for conventional deep neural networks (DNNs) to directly extract meaningful local features [ali2018yolo3d, yang2018pixor, choy20194d]. To overcome this limitation, point-based neural networks (PNNs) [he2025pointrwkv, deng2023pointvector, qian2022pointnext, qi2017pointnet++, qi2017pointnet, phan2018dgcnn, zhao2019pointweb, wang2019dynamic] incorporate specialized point operations, such as sampling, neighbor searching, and gathering, to effectively extract learnable local features from unordered point cloud. These tailored operations significantly improve the performance of PNNs in 3D perception tasks. However, these operations often involve irregular memory access patterns and intensive computation, which pose significant challenges to the efficiency of PNNs on edge hardware.
@@
However, as PNN evolves, the scale of point cloud data has expanded from 1K to over 200K in benchmarks [sun2022modelnet40, armeni20163d], with modern LiDAR sensors producing 30K–300K points per frame in real life [roriz2024survey, li2023instant]. This increase introduces new challenges, with point operations emerging as the primary bottleneck, as shown in Fig. 4. Conventional point operations require iterative global memory access and all-to-all computation, referred to as global search. The iterative nature and data dependency severely limit decomposition and parallelization, both essential for efficient large-scale processing. Partitioning can mitigate global search by localizing computations and memory accesses. However, prior works mainly target small-scale inputs (<4K<4K points) and overlook this parallelization issue. As a result, their hardware designs are misaligned with the demands of modern workloads, causing significant performance degradation on large-scale point clouds.
@@ VII Conclusion
Acknowledgements
This work was supported in part by NSF grants 2328805 and 2112562, by ARO W911NF-23-2-0224, as well as the Beyond the Horizon Initiative through the Pratt School of Engineering, Duke University. The authors would like to thank the anonymous reviewers for their constructive feedback, which helped improve the quality of this work. We also extend our gratitude to Junyao Zhang, Haoxuan Shan, Jonathan Ku for their technical support and insightful discussions.

### v2

@@ Abstract
Three-dimensional (3D) point clouds are increasingly used in applications such as autonomous driving, robotics, and virtual reality (VR). Point-based neural networks (PNNs) have demonstrated strong performance in point cloud analysis, originally targeting small-scale inputs. However, as PNNs evolve to process large-scale point clouds with hundreds of thousands of points, all-to-all computation and global memory access in point cloud processing introduce substantial overhead, causing $O(n^{2})$ computational complexity and memory traffic where $n$ is the number of points. Existing accelerators, primarily optimized for small-scale workloads, overlook this challenge and scale poorly due to inefficient partitioning and non-parallel architectures. To address these issues, we propose FractalCloud, a fractal-inspired hardware architecture for efficient large-scale 3D point cloud processing. FractalCloud introduces two key optimizations: (1) a co-designed Fractal method for shape-aware and hardware-friendly partitioning, and (2) block-parallel point operations that decompose and parallelize all point operations. A dedicated hardware design with on-chip fractal and flexible parallelism further enables fully parallel processing within limited memory resources. Implemented in 28 nm technology as a chip layout with a core area of 1.5 $mm^{2}$, FractalCloud achieves 21.7× speedup and 27× energy reduction over state-of-the-art accelerators while maintaining network accuracy, demonstrating its scalability and efficiency for PNN inference. The code for FractalCloud is available at https://github.com/Yuzhe-Fu/FractalCloud.
11footnotetext: Corresponding Author: Changchun Zhou and Cong Guo
@@ I Introduction
With the rapid advancement of 3D sensing technologies [rani2024advancements, guo2024revisiting, yan2024spot, zhou2023sagitta], such as LiDAR [roriz2024survey] and stereo cameras [dong20231920, dong20214, dong2021139], point cloud has become prevalent in a wide range of edge applications, including autonomous driving [liu2025towards, cui2021deep], robotics [kastner20203d], augmented reality [wang2018point], and virtual reality [stets2017visualization, li2024fusion]. The points in a 3D point cloud are inherently unordered and irregularly distributed, making it challenging for conventional deep neural networks (DNNs) to directly extract meaningful local features [ali2018yolo3d, yang2018pixor, zhang2023pids, choy20194d]. To overcome this limitation, point-based neural networks (PNNs) [he2025pointrwkv, deng2023pointvector, qian2022pointnext, qi2017pointnet++, qi2017pointnet, phan2018dgcnn, zhao2019pointweb, wang2019dynamic] incorporate specialized point operations, such as sampling, neighbor searching, and gathering, to effectively extract learnable local features from unordered point cloud. These tailored operations significantly improve the performance of PNNs in 3D perception tasks. However, these operations often involve irregular memory access patterns and intensive computation, which pose significant challenges to the efficiency of PNNs on edge hardware.
@@
However, as PNN evolves, the scale of point cloud data has expanded from 1K to over 200K in benchmarks [sun2022modelnet40, armeni20163d], with modern LiDAR sensors producing 30K–300K points per frame in real life [roriz2024survey, li2023instant]. This increase introduces new challenges, with point operations emerging as the primary bottleneck, as shown in Fig. 4. Conventional point operations require iterative global memory access and all-to-all computation, referred to as global search. The iterative nature and data dependency severely limit decomposition and parallelization, both essential for efficient large-scale processing [guo2020accelerating, guo2024accelerating]. Partitioning mitigates global search by localizing computations and memory accesses. However, prior works mainly target small-scale inputs (<4K<4K points) and overlook this parallelization issue. As a result, their hardware designs are misaligned with the demands of modern workloads, causing significant performance degradation on large-scale point clouds.
@@ VII Conclusion
Acknowledgments
This work was supported in part by NSF grants 2328805 and 2112562, by ARO W911NF-23-2-0224, as well as the Beyond the Horizon Initiative through the Pratt School of Engineering, Duke University. The authors would like to thank the anonymous reviewers for their constructive feedback, which helped improve the quality of this work. We also extend our gratitude to Junyao Zhang, Haoxuan Shan, Mingyuan Ma, and Jonathan Ku for their technical support and insightful discussions.
Appendix A Artifact Appendix
A-A Abstract
Our artifact contains the algorithmic implementation of FractalCloud, enabling the evaluation of network performance under the proposed fractal-inspired point cloud processing architecture. The artifact includes the PyTorch implementation of the Fractal operation and Block-Parallel Point Operations, scripts to reproduce classification and segmentation results on ModelNet40 and S3DIS, and pretrained models for PointNet++, PointNeXt-S, and PointVector-L. The purpose of this artifact is to validate algorithmic correctness and network performance as reported in the paper.
A-B Artifact check-list (meta-information)
  • Algorithm: Fractal operation; block-parallel point operations; Point cloud neural networks
  • Program: Python (PyTorch implementation)
  • Compilation: N/A (Conda environment and Docker)
  • Transformations: N/A
  • Binary: Docker Image
  • Model: PointNet++, PointNeXt-S, PointVector-L
  • Data set: ModelNet40, S3DIS
  • Run-time environment: Ubuntu 24.04.3 LTS (Host), Docker container with Ubuntu 16.04.7 LTS, Python 3.7.16, PyTorch 1.10.1, CUDA 11.3
  • Hardware: x86-64 server with one NVIDIA TITAN RTX (24GB)
  • Run-time state: Standard inference with pretrained models
  • Execution: Python scripts
  • Metrics: Classification accuracy, segmentation mIoU
  • Output: Reported accuracy and segmentation metrics
  • Experiments: Classification on ModelNet40; segmentation on S3DIS
  • How much disk space required (approximately)?: $\sim$96GB total (including Docker, dependencies, datasets, pretrained models, and cache)
  • How much time is needed to prepare workflow (approximately)?: $\sim$20 minutes (Docker)
  • How much time is needed to complete experiments (approximately)?: 5–10 minutes for classification; 2–14 hours for segmentation
  • Publicly available?: Yes (https://github.com/Yuzhe-Fu/FractalCloud)
  • Code licenses (if publicly available)?: MIT license
  • Data licenses (if publicly available)?: Original dataset licenses
  • Workflow automation framework used?: N/A
  • Archived (provide DOI)?: https://zenodo.org/records/17851585
A-C Description
A-C1 How to access
The source code is publicly available at: https://github.com/Yuzhe-Fu/FractalCloud. The artifact sources are also archived at Zenodo: https://zenodo.org/records/17851585. Pretrained weights for all evaluated models (PointNet++, PointNeXt-S, and PointVector-L) and inference logs for evaluated workflows are available at: https://huggingface.co/YuzheFu/FractalCloud.
A-C2 Hardware dependencies
A server equipped with a GPU is required for evaluation. The artifact has been tested with an NVIDIA TITAN RTX GPU (24GB).
A-C3 Software dependencies
The artifact relies on the following software components:
  • Host OS: Ubuntu 24.04.3 LTS
  • Docker (optional): Docker 28.3.3
  • Docker Image (provided): Ubuntu 16.04.7 LTS with all required dependencies preinstalled
  • Python 3.7.16
  • Pytorch 1.10.1
  • CUDA 11.x (only required for local installation without the provided Docker image)
  • Conda 4.9.2
A-C4 Data sets
The evaluated datasets are:
  • ModelNet40 (classification)
  • S3DIS (segmentation)
A-C5 Models
  • PointNet++
  • PointNeXt-S
  • PointVector-L
Pretrained weights and inference logs (exact outputs reported in the paper) are available at https://huggingface.co/YuzheFu/FractalCloud for validating reproduced results.
A-D Installation
Installation can be done through the provided Docker image or via running local setup scripts. A detailed step-by-step installation guide is provided in the project README file.
A-E Experiment workflow
Example execution commands to reproduce classification and segmentation results are included in the README. Running the provided scripts will produce network accuracies matching the results reported in the paper.
A-F Evaluation and expected results
The artifact is expected to reproduce the network accuracy metrics reported in the paper. Specifically, it validates the numerical consistency of the fractal implementation and its impact on network performance across multiple point cloud models.
A-G Experiment customization
Users can modify experiment configurations, model architectures, and dataset paths by editing the script arguments and YAML configuration files as described in the project README.
A-H Notes
N/A
A-I Methodology
Submission, reviewing and badging methodology: